2.5 g of Al was made to react with 150 mL 0.25 M HCl. What was the limiting reactant? How many g AlCl3 was formed? How many L of H was collected if the reaction took place at 35 deg C and 720 mm Hg? Make the appropriate chemical reaction. balance the chemical equation first. 


1
Expert's answer
2021-02-17T07:41:36-0500

The balanced chemical equation is-

"2Al+6HCl \rightarrow 2AlCl_3+3H_2"


Mass of Al= 2.5g


Moles of Al = "\dfrac{2.5}{2.7}=0.09259"


Volume of "HCl=150ml"

Molarity of "HCl=0.25M"


No. of moles of "HCl= \dfrac{(150\times 0.25)}{1000} =0.0375"


Here the limiting reagent is Al


So, the number of moles of "AlCl_3=\dfrac{2}{2}\times 0.0925=0.0925"


Mass of "AlCl_3= Moles \times \text{Molar mass of } AlCl_3=0.0925\times133.5=12.41g"


Also, Pressure "P = 720mmHg" = 1atm


Temprature, "T= 35^{\circ} = 35+273=308K"


Moles of "H_2 = \dfrac{(3\times 0.093)}{2}=\dfrac{0.279}{2}=0.1395"


Using the equation "PV=nRT"


"1\times V=6.1395\times 8.314\times 306"


"V=\dfrac{3957.21}{1000}=0.35721=0.36L"


Hence, the volume of "H_2 \text{ is } 0.36L"